During a science experiment, Jean created a 36% alcohol solution by mixing 11 fluid ounces of a 50% alcohol solution with a certain amount of a 14$ alcohol solution. How many fluid ounces of the 36% alcohol solution did she create?
A. 25 fl oz
B. 18 fl oz
C. 20 fl oz
D. 7 fl oz

Answer by richwmiller(17219)

Let C denote the 14% solution and D denote the final 36% solution.
We have 11 oz of 50% solution.
Using the method of alligation.
14= 36 - 50
22= 14 - 36
E.g. There are 22 parts of C for every 14 parts of D
then there are 11 parts of z, since (22/14)=11/z
x= 7.0= 11/(22/14) oz of 14% solution to add to your 11 oz of 50% solution.
y= 18.0= (11+7.0) oz of the final 36% solution.
check
0.14 * 7.0 + 0.5 * 11 = 0.36*(11+7.0)
0.98 + 5.5 = 0.36*(18.0)
6.48 = 6.48
ok
50         22        11 oz of 50% solution
      36            18 oz of 36% solution
14         14 14/22*11=7 oz of 14% solution
           36